OpenStudy (anonymous):

What is the slope of the line x = −5? (1 point) Undefined 0 −5 5 Anyone know please?

Nnesha (nnesha):

vertical line = 0 slope (remember slope = rise /over) horizontal line = undefined

OpenStudy (anonymous):

So it's 0?

Nnesha (nnesha):

yes right bec = -5 would be the vertical line

OpenStudy (anonymous):

Ok ty

OpenStudy (anonymous):

It said it was wrong :(

Nnesha (nnesha):

oh god facepalm i switched x and y i'm soo sorry

Nnesha (nnesha):

x=vertical line = undefined slope y=horizontal line = 0 slope bec x values would be the same and if you use the formula \[\huge\rm \frac{ y_2-y_1 }{ x_2 -x_1 }\] we would get 0 at the denominator (bec the line have same x-coordinate ) and when you subtract same two numbers you will get and if we get 0 at the denominator then answer would be undefined

Nnesha (nnesha):

 same x value for ally-coordinates

Nnesha (nnesha):

so \[\huge\rm \frac{ y_2 - y_1 }{ x_2-x_1 }= \frac{ y_2 -y_1}{ -5-(-5)}\]

Nnesha (nnesha):

did you get it ? and i'm sorry ...:(

OpenStudy (anonymous):

its ok :( and i got a C- atleast not a D :/